OpenStudy (anonymous):

If 25 power y = 8 power x, and 125 power y = 4 power (x-1), Find "x".

OpenStudy (anonymous):

are you saying 25^8^x and 125^4^(x-1)?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

25 power y. what does that mean then if it doesn't mean to the power of (^)

OpenStudy (anonymous):

is it 25^y = 8^x and 125^y = 4^(x-1)

OpenStudy (anonymous):

\[25^{y} = 8^{x}\]\[yln25 = xln8\]\[y = \frac{x \ln8}{\ln25}\] and \[125^{y} = 4^{x-1}\]\[y \ln125 = (x-1) \ln4\]\[y = \frac{(x-1)\ln4}{\ln125}\] Substitute for y so the two equations are equal:\[\frac{x \ln8}{\ln25} = \frac{(x-1)\ln4}{\ln125}\]Then solve for x:\[x \ln8 \ln125 = (x-1)\ln4 \ln25\]\[x \ln8 \ln125 = x \ln4 \ln25 -\ln4 \ln25 \]\[ (\ln8 \ln125 - \ln4 \ln25)x = -\ln4 \ln25 \]\[x = \frac{ -\ln4 \ln25 }{ (\ln8 \ln125 - \ln4 \ln25)}\]\[x = \ln4 = 1.38629\]

OpenStudy (zarkon):

\[25^y=8^x\Rightarrow 5^{2y}=2^{3x}\] \[125^y=4^{x-1}\Rightarrow5^{3y}=2^{2(x-1)}\] \[(5^{2y})^{3/2}=(2^{3x})^{3/2}\Rightarrow5^{3y}=2^{9x/2}\] \[2^{2(x-1)}=2^{9x/2}\Rightarrow2(x-1)=9x/2\Rightarrow4x-4=9x\Rightarrow x=\frac{-4}{5}\]